2. Find r(t) if r(t) = 5tại + vtj ifr(0) = 2i + j.
Answer #1

solution r (t) = 8+27 +ve ¢ (0) = i + ن و ) ا= (t) 8 آن هم بلرc + و - ری 1 + ( . (ه) 8 G) - + 2 q given ro) + ى comparing eq